OpenStudy (anonymous):

how do you solve 5/8 under the radical * 2/6 under the radical?

OpenStudy (diyadiya):

\[\sqrt{\frac{5}{8}} \times \sqrt{\frac{2}{6}}\]??

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

\[\sqrt{\frac{5}{8} \times \frac{2}{6}}=\sqrt{\frac{10}{48}}=\sqrt{\frac{5}{24}}\]

OpenStudy (anonymous):

thank you